OpenStudy (anonymous):

When 25 2/3 is added to 19 3/4 the answer is 44 5/12. A. True B. False

OpenStudy (anonymous):

How is it A?

OpenStudy (nightmarenight):

wait no it's B

OpenStudy (anonymous):

Explanation please.

OpenStudy (nightmarenight):

get a common denominator: 12

OpenStudy (nightmarenight):

so now you have 25 and 8/12 and 19 and 9/12

OpenStudy (nightmarenight):

since 8/12 + 9/12 = 17/12 you take 12 and it goes into 17 once, so now you have 1 5/12

OpenStudy (nightmarenight):

add it all up and you get 45 5/12

OpenStudy (nightmarenight):

let me look it over to make sure I got the right answer

OpenStudy (nightmarenight):

do you need to show your work? I can put it up here

OpenStudy (nightmarenight):

\[25\frac{ 2 }{ 3 } + 19\frac{ 3 }{ 4 } \]

OpenStudy (nightmarenight):

take the fractions and find a common denominator

OpenStudy (nightmarenight):

\[\frac{ 2 }{ 3 } + \frac{ 3 }{ 4 } \]

OpenStudy (nightmarenight):

so do 3 x 4

OpenStudy (nightmarenight):

\[\frac{ 2 }{ 3 } x \frac{ 4 }{ 4 } = \frac{ 8 }{ 12 }\]

OpenStudy (nightmarenight):

\[\frac{ 3 }{ 4 } x \frac{ 3 }{ 3 } = \frac{ 9 }{ 12 }\]

OpenStudy (nightmarenight):

\[now, \frac{ 8 }{ 12 } + \frac{ 9 }{ 12 } = \frac{ 17 }{ 12 }\]

OpenStudy (nightmarenight):

since \[\frac{ 17 }{ 12 }\] is not a correct fraction, you want to take 12 and it goes into 17 once, now you have \[1\frac{ 5 }{ 12 }\]

OpenStudy (nightmarenight):

You're going to take your whole numbers from before and add them to \[1\frac{ 5 }{ 12 }\]

OpenStudy (nightmarenight):

\[1\frac{ 5 }{ 12 } + 25 + 19= 45 \frac{ 5 }{ 12 }\]

OpenStudy (nightmarenight):

Did you understand that?